"Top Gear" magazine published an article stating that car brands are purchased in equal numbers. To examine this claim, a sample of cars on a major highway was collected and the numbers of 5 major car brands was reported. The numbers are given in the table. Does this sample contradict the magazine's claim at the 0.10 level of significance.  

Honda Toyota Ford   GM   Hyundai
121 105 98 92 102
Answer #1

Answer: Top Gear" magazine published an article stating that car brands are purchased in equal numbers. To examine this claim, a sample of cars on a major highway was collected and the numbers of 5 major car brands was reported. The numbers are given in the table. Does this sample contradict the magazine's claim at the 0.10 level of significance.  

The hypothesis test:

Ho: The car brands are purchased in equal numbers = 1/5

Ha: Some of the car brands purchased is not in equal numbers ≠ 1/5

n = 121 + 105 + 98 + 92 + 102 = 518

Test statistic χ2 = (Observed value - Expected value)^2/ Expected value

Expected value = np

Expected value = 518 * 1/5

Expected value = 103.6

Brands Observed (O) Expected (E) (O-E)^2/E
Honda 121 103.6 2.9224
Toyota 105 103.6 0.0189
Ford 98 103.6 0.3027
GM 92 103.6 1.2988
Hyundai 102 103.6 0.0247
Total 518 518 4.5675

Test statistic χ2 = 4.5675

P-value:

df = (r-1)(c-1)

df = (5-1)(2-1)

df = 4

P-value = 0.3347

Since, P-value (0.33) > α (0.05) significance level.

We do not reject the null hypothesis Ho.

Conclusion:

Do not reject Ho. There is not sufficient evidence to conclude that some of the car brands purchased is not in equal numbers.